OpenStudy (anonymous):

The windows of a downtown office building are arranged so that each floor has 6 fewer windows than the floor below it . If the ground floor has 52 windows, how many windows are on the 8th floor ? A. 10 B. 8 C. 4 D. 6

OpenStudy (anonymous):

What do you think?? How would you frame the equation??

OpenStudy (anonymous):

1st floor-52 windows how many windows on the 2nd floor?

OpenStudy (anonymous):

46 ?

OpenStudy (anonymous):

Yes on 2nd floor , how many??

OpenStudy (anonymous):

46 windows ?

OpenStudy (anonymous):

s_b and I are trying to help. we could give you the answer but that wouldn't really help you to learn

OpenStudy (anonymous):

Sorry I meant the 3 rd floor how much??

OpenStudy (anonymous):

i know ?

OpenStudy (anonymous):

40 windows on the 3rd floor ?

OpenStudy (anonymous):

@thomasj ground floor has 52 windows, not first floor :P

OpenStudy (anonymous):

sorry. you are right.

OpenStudy (anonymous):

but wouldnt it be 64 windows on the 3rd floor since is says the ground floor has 52 windows and it says that every window has 6 fewer than the one below it,?

OpenStudy (anonymous):

so you would coubt up since there arent any floors below ground floor?

OpenStudy (anonymous):

See you are saying. I had 6 chocolates earlier. Now I have 1 chocolate lesser , how many chocolates do I have now?? You are saying 7 . which is wrong.

OpenStudy (anonymous):

so you would subtract from 52?

OpenStudy (anonymous):

Yes exactly. Now tell me how do you get 46 windows. Show me subtraction procedure. Also for 2nd floor

OpenStudy (anonymous):

52-6 = 46 windows on 1st floor

OpenStudy (campbell_st):

use the sum of an arithmetic series a = 52 n = 9 and d = 6

OpenStudy (anonymous):

On 2nd floor ??

OpenStudy (anonymous):

40 on second floor

OpenStudy (anonymous):

I mean the subtraction procedure..

OpenStudy (anonymous):

46-6 = 40

OpenStudy (anonymous):

Or 52- 2(6) = 40

OpenStudy (anonymous):

So basically it is an Arithmetic progression as @campbell_st said which is having a=52 , d = 6. Now what would be n???

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

Sorry it is d = -6.

OpenStudy (campbell_st):

use the term in an arithmetic\[T _{8}= 52 + (*-1)\times-6\] series where a = 52, n = 8\[t _{8} = 52 - (8 -1)\times -6\] and d= -6

OpenStudy (anonymous):

why -6 ? is it because your subtracting ?

OpenStudy (anonymous):

@CCarmichael18 , yes you are right. Now use the formula I told you in the previous sum and tell me the answer

OpenStudy (anonymous):

the answer is 10 ?

OpenStudy (anonymous):

a8= 10 windows

OpenStudy (campbell_st):

oops n = 9.... its G to 8 = 9

OpenStudy (anonymous):

there are 10 windows on floor 8

OpenStudy (anonymous):

No there are 4 windows. n= 8

OpenStudy (anonymous):

oh okay, thank you

OpenStudy (anonymous):

Welcome :) and thanks to @campbell_st and @thomasj

OpenStudy (anonymous):

the answer is 10.. 100% sure , just took the quiz , use the formula a8=52(8-1)= -6 a8=52-42 a8=10 10 windows are on the eighth floor .. message me if anyone need help :)

OpenStudy (anonymous):

THE ANSWER IS "10". HERES THE WORK BELOW... lol it took me a while to do this :)